# When Large Language Models Meet Vector Databases: A Survey

### v1

@@ 4.0.3 Data preprocessing
While text embedding is considered efficient for processing long text and its retrieval, it comes with a dazzling challenge for building every text-based knowledge database. It is observed that vector retrieval would have fundamentally better performance when precise chunks of text with clear meaning are applied, this raises an interesting question: how to have a proper (probably unified) embedding methodology for every raw text?
On the other hand, dimension reduction for vector data is also considered important due to the curse of dimensionality theory. The current vector data dimension may be not cost-effective, as vectors are large: e.g., a 1536 dimensional vector is about 6 kilobytes, with scaling up to 1 billion, the data size would be 6 terabytes. This requires a more efficient vector data dimension reduction algorithm that matches vector databases and search algorithms.
@@ 4.0.4 Knowledge conflict
Conflicting knowledge from the same or different knowledge bases presents a non-trivial challenge for both humans and LLMs to distinguish the correct piece of knowledge. This conflict becomes particularly pronounced when integrating multiple knowledge bases, each potentially carrying its own biases or inaccuracies. Resolving such conflicts requires robust conflict resolution strategies that can assess the reliability of sources and the context of data to determine the most accurate information.
4.0.5 Data management systems for LLMs with multi-tenancy
This challenge is particularly in maintaining the isolation and security of distinct tenants’ data while ensuring efficiency, where isolation is crucial for privacy and security, yet it must be balanced with the need for resource sharing to ensure cost-effectiveness.
@@ 5 Conclusion
In this paper, we present a systematic review of recent advances in combinations of LLMs and vector databases. We also introduce recent VecDB+LLMs applications with distinct prototypes that categorize existing works from various perspectives and interdisciplinary studies. Our study also demonstrates both the research and engineering challenges in this fast-growing field and suggests directions for future work.

### v2

@@ 5.0.3 Data preprocessing
While text embedding is considered efficient for processing long text and its retrieval, it comes with a dazzling challenge for building every text-based knowledge database. It is observed that vector retrieval would have fundamentally better performance when precise chunks of text with clear meaning are applied. This raises an interesting question: how do we have a proper (probably unified) embedding methodology for every raw text?
On the other hand, dimension reduction for vector data is considered important since compressing high-dimensional vectors will use less memory and disk resources. From this view, the current vector data dimension is not cost-effective, as the vectors are large, e.g., a 1536-dimensional vector embedded only by OpenAI’s lightweight model text-embedding-3-small  [^6] https://platform.openai.com/docs/guides/embeddings is about 6 kilobytes; with scaling up to 1 billion, the data size would be 6 terabytes. Also with the curse of dimensionality theory, data sparsity increases with dimensionality, resulting in hardship in distinguishing meaningful similarities. Moreover, current dimension reduction algorithms like PQ contribute to a memory-efficient vector data representation, whereas lossy data compression still hurts the recall quality Chen et al. (2021). This requires a more efficient vector data dimension reduction algorithm for VecDBs and searches.
5.0.4 Data management systems for LLMs with multi-tenancy
Both for VecDBs and LLMs, this challenge is particularly in preventing data leaks between tenants and maintaining the isolation and security of distinct tenants’ data while ensuring efficiency, where isolation is crucial for privacy and security. On the other hand, it must be balanced with the need for dynamic resource sharing to ensure resources are flexibly allocated and deallocated.
5.0.5 Cost-effective and scalable data storage and retrieval
While state-of-the-art vector search algorithms are mainly memory-based, memory resources in the cloud are considered expensive and volatile. With billion-scale amounts of data, naively handling them in memory resources is bankrupting. On the other hand, disk-based search methods have downgraded performance that does not satisfy real-time applications with low latency tolerance. This calls for a more cost-effective way to ensure the performance of LLM applications.
@@ 5.0.6 Knowledge conflict
Conflicting knowledge from the same or different knowledge bases presents a non-trivial challenge for both humans and LLMs to distinguish the correct piece of knowledge. This conflict becomes particularly prominent when integrating multiple knowledge bases where each potentially carries its own biases or inaccuracies, such as when different datasets provide contradictory information about the same fact or when biases inherent in data sources lead to inconsistent conclusions. While VecDBs cannot naturally distinguish, resolving such conflicts requires robust conflict resolution strategies that can assess the reliability of sources and the context of data to determine the most accurate information.
@@ 6 Conclusion
In this paper, we present a systematic review of recent advances in combinations of LLMs and VecDBs. We summarize the rationales for viewing LLMs with VecDBs and also introduce certain applications that combine LLMs and VecDBs with distinct prototypes that categorize existing works from various perspectives and interdisciplinary studies. Our study also demonstrates both the research and engineering challenges in this fast-growing field and suggests directions for future work.